What is the H^+ ion concentration in a 2.1x10^-4 M Ca(OH)2 solution? Please explain in detail and do step by step. Thanks!
Answer #1

1Ca(OH)2 >Ca*2+20H 1mol of Ca(OH)2 producing 2 moles of OH Concentration of Ca(OH)2 2.1×10-4M Concentration of OH- 2x (2.1x10
Answer #2

Ca(OH)2  \rightarrow    Ca2+ + 2OH-

Here, we can see that one mole of Ca(OH)2 produces two moles of OH-

Therefore, concentration of OH- = [OH-] = 2 x (2.1x10-4 M ) = 4.2 x 10-4 M

We have, [H+][OH-] = Kw (Kw = 1 x 10-14 )

Therefore, [H+] = (1 x 10-14) /(4.2 x 10-4 M) = 2.3809 x 10-11 M

H+ ion concentration in a 2.1x10-4 M Ca(OH)2 solution = 2.3809 x 10-11 M